1) Allow f(x) = x3, and g(x) = 2x. Determine f(g(1)).

2) Enter the formula for a sine function with an amplitude of 5, a period of 90 degrees, a shift of 45 degrees to the right and 3 units upwards.

Determine the exact value of the expression sin (v) + cos (v) given that tan (v) = 7/3.

3) Solve the equations in the range 180≤x≤360 degrees. Answer with one decimal.

a) tan (x) = 3.2 (I would like a full solution for this, please)

b) sin (x) = 1.6
Answer #1

( f(3) = ??, gº =23 fCq(1) = f(20)) = f(2)= (2)3=5 Amplitude (A) = 5 90 (2) c) period (360)= B: 360 90 4. place shift ) = 45°3 180 <a 5360 (9) tany)=312 = 32 To S 16 5 X tan (16 used desmes, o/ 252.646 online graph Matter (5) sin(= 116 a = sino (1:6)